OpenStudy (anonymous):

How do you find the volume of a square based pyramid with 8cm sides and slopping edges 10cm in length?

OpenStudy (cwrw238):

volume of a square pyramid = (1/3)b^2 h where b = length of side of base and h = altitude you can find the altitude by using pythagoras theorem

OpenStudy (cwrw238):



OpenStudy (cwrw238):

h^2 = 10^2 - 4^2

OpenStudy (cwrw238):

do you get it?